OpenStudy (anonymous):

A combination lock should probably be called a permutation lock, because the order of the numbers does matter. How many different 3-number permutations are there on a lock with 36 digits?

OpenStudy (mathstudent55):

You need: \(_nP_r = \dfrac{n!}{ (n - r)! } \) \( _{36}P_3 = \dfrac{36!}{ (36 - 3)! } \)